A spider has 8 legs and no wings. A dragonfly has 6 legs and 2 pairs of wings. A cicada has 6 legs and one pair of wings. We have all 3 kinds and a total of 18 insects in a cage. This gives us a total of 118 legs and 20 pairs of wings. How many insects do we have of each kind?
Answer by checkley71(8403)

8L=S
6L+2W=D
6L+1W=C
S+D+C=18
TO GET 118 LEGS & 20 PAIR OF WINGS WE NEED
5 SPIDERS+7 DRAGON FLIES+6 CICADA
5 SPIDERS HAVE 40 LEGS
7 DRAGONS HAVE 42 LEGS & 14 SETS OF WINGS
6 CICADA HAVE 36 LEGA & 6 SETS OF WINGS
------------------------------------------------------------
18 INSECTS HAVE 118 LEGS & 20 SETS OF WINGS (MAKE UP A TABLE WITH ALL THE DATA AND START FILLING IN THE BOXES)